How many ounces of 22% acid solution must be added to 10 ounces of 8% solution to obtain 12% solution?
Answer by richwmiller(17219)

We have 10 oz of 8% solution
We want to add some 22% solution to make 12% solution
0.22*x+0.08*10=0.12(10+x)
0.22*x+0.8=1.2+0.12x
0.22x-0.12x=1.2-0.8
0.1x=0.4
x=4 oz at 22% (0.88) is added to 10 oz at 8%(0.8) to make 12% of 14 oz(1.68)
check
0.22*4+0.08*10=0.12(10+4)
0.88+0.8=0.12(14)
1.68=1.68
ok
codemixadd